# TutorLLM: Customizing Learning Recommendations with Knowledge Tracing and Retrieval-Augmented Generation

### v1

@@ 1. Introduction
To bridge the gap between the adaptability of LLMs and the personalized approach of educational recommender systems, we propose a novel framework: the Personalized Educational Recommender LLM System (TutorLLM), based on Knowledge Tracing (KT) and Retrieval-Augmented Generation (RAG). To the best of our knowledge, we are the first to incorporate KT technology into an LLM to achieve a personalized recommendation learning framework. The TutorLLM comprises three integral components. The first component is a KT model, which is developed based on the Multi-Features with Latent Relations BERT Knowledge Tracing (MLFBK) model (Li et al., 2023). This component not only gathers data on student interactions and performance but also assimilates information from dialogues with LLMs, offering insights into student capabilities, learning states, and the complexities of the knowledge being acquired. The second is a Scraper. This component collects text content during online learning sessions to provide context-specific background knowledge that enhances the relevance and accuracy of the LLM’s responses. The third component is an RAG Enhanced LLM. This component, utilizing the GPT-4 API, integrates inputs from both the Scraper and the KT module to deliver precise, personalized responses and learning content recommendations. Additionally, for even more tailored interaction, students can manually upload learning materials. Figure 1 shows the overall model of the architecture. The text content provided by Scraper mitigates hallucinations by dynamically retrieving and incorporating context-specific knowledge from relevant course materials, ensuring accurate and reliable information. Additionally, the integration of knowledge tracing allows TutorLLM to deliver highly personalized recommendations and responses, tailored to each student’s learning progress and needs.
@@ 2. Related Work
Educational Recommender Systems provide personalized learning experiences by tailoring educational content based on students’ needs, preferences, and progress (Ricci et al., 2011). Traditional systems use collaborative filtering, content-based filtering, or hybrid approaches to suggest relevant resources. Combining these systems with AI technologies like LLMs and KT models can significantly enhance personalization and performance in learning outcomes (Chen et al., 2020; Zhang et al., 2019).
Knowledge Tracing monitors and predicts students’ knowledge states over time by tracking interactions with learning materials and assessments (Corbett and Anderson, 1994; Piech et al., 2015). It enables personalized recommendations and targeted interventions to address the knowledge gaps of the students. Traditional methods like Bayesian Knowledge Tracing (BKT) (Pelánek, 2017) and Deep Knowledge Tracing (DKT) (Piech et al., 2015) have been widely used in Educational Recommender Systems, leveraging historical data to predict future performance and learning needs (Liu et al., 2021; Pandey and Pandey, 2019). Recent models, such as Multi-Features with Latent Relations BERT Knowledge Tracing (MLFBK), enhance the accuracy and depth of predictions (Chen et al., 2023; Lee et al., 2022; Tiana et al., 2021; Li et al., 2023).
@@ 3. Implementation of TutorLLM
the UI if the TutorLLM
Knowledge Tracing Model: For the knowledge tracing part, we use MLFBK(Li et al., 2023) to incorporate multi-feature embedding and latent Relations to capture students’ learning state. MLFBK consists of three parts: embedding, BERT-based architecture, and correctness sequence output.
@@ 4. User Study
Participants & Study Setup. To evaluate the effectiveness of the TutorLLM educational tool on student learning outcomes, 30 first-year undergraduate students from XXX University (anonymous for review) enrolled in a linear algebra module were randomly allocated to one of three groups in a controlled experiment. The control group exclusively used general LLMs (the general web version of chatGPT4) for the duration of the study. Experimental Group 1 was exposed to general LLMs during the first week, followed by TutorLLM in the second week, whereas Experimental Group 2 utilized TutorLLM throughout both weeks. The key objective was to determine if integrating TutorLLM could enhance learning performance compared to the general LLMs.

### v2

@@ 1. Introduction
To bridge the gap between the adaptability of LLMs and the personalized approach of educational recommender systems, we propose a novel framework: the Personalized Educational Recommender LLM System (TutorLLM), based on Knowledge Tracing (KT) and Retrieval-Augmented Generation (RAG). To the best of our knowledge, we are the first to incorporate KT technology into an LLM to achieve a personalized recommendation learning framework. The TutorLLM comprises three integral components. The first component is a KT model, which is developed based on the Multi-Features with Latent Relations BERT Knowledge Tracing (MLFBK) model (Li et al., 2023a). This component not only gathers data on student interactions and performance but also assimilates information from dialogues with LLMs, offering insights into student capabilities, learning states, and the complexities of the knowledge being acquired. The second is a Scraper. This component collects text content during online learning sessions to provide context-specific background knowledge that enhances the relevance and accuracy of the LLM’s responses. The third component is an RAG Enhanced LLM. This component, utilizing the GPT-4 API, integrates inputs from both the Scraper and the KT module to deliver precise, personalized responses and learning content recommendations. Additionally, for even more tailored interaction, students can manually upload learning materials. Figure 1 shows the overall model of the architecture. The text content provided by Scraper mitigates hallucinations by dynamically retrieving and incorporating context-specific knowledge from relevant course materials, ensuring accurate and reliable information. Additionally, the integration of knowledge tracing allows TutorLLM to deliver highly personalized recommendations and responses, tailored to each student’s learning progress and needs.
@@ 2. Related Work
Educational Recommender Systems provide personalized learning experiences by tailoring educational content based on students’ needs, preferences, and progress (Ricci et al., 2011). Traditional systems use collaborative filtering, content-based filtering, or hybrid approaches to suggest relevant resources. Combining these systems with AI technologies like LLMs and KT models can significantly enhance personalization and performance in learning outcomes (Chen et al., 2020; Zhang et al., 2019; Wang et al., 2024a, 2023, c, b).
Knowledge Tracing monitors and predicts students’ knowledge states over time by tracking interactions with learning materials and assessments (Corbett and Anderson, 1994; Piech et al., 2015). It enables personalized recommendations and targeted interventions to address the knowledge gaps of the students. Traditional methods like Bayesian Knowledge Tracing (BKT) (Pelánek, 2017) and Deep Knowledge Tracing (DKT) (Piech et al., 2015) have been widely used in Educational Recommender Systems, leveraging historical data to predict future performance and learning needs (Liu et al., 2021; Pandey and Pandey, 2019). Recent models, such as Multi-Features with Latent Relations BERT Knowledge Tracing (MLFBK), enhance the accuracy and depth of predictions (Li et al., 2021; Chen et al., 2023; Lee et al., 2022; Tiana et al., 2021; Li et al., 2023a; Li, 2023; Li et al., 2023c, 2024b, b, 2024a).
@@ 3. Implementation of TutorLLM
the UI of the TutorLLM
Knowledge Tracing Model: For the knowledge tracing part, we use MLFBK(Li et al., 2023a) to incorporate multi-feature embedding and latent Relations to capture students’ learning state. MLFBK consists of three parts: embedding, BERT-based architecture, and correctness sequence output.
@@ 4. User Study
Participants & Study Setup. To evaluate the effectiveness of the TutorLLM educational tool on student learning outcomes, 30 first-year undergraduate students from the University of Southampton enrolled in a linear algebra module were randomly allocated to one of three groups in a controlled experiment. The control group exclusively used general LLMs (the general web version of chatGPT4) for the duration of the study. Experimental Group 1 was exposed to general LLMs during the first week, followed by TutorLLM in the second week, whereas Experimental Group 2 utilized TutorLLM throughout both weeks. The key objective was to determine if integrating TutorLLM could enhance learning performance compared to the general LLMs.